OpenStudy (anonymous):

Find the angle of intersection of the plane -5x-2y+4z=2 with the plane 3x+3y+5z=5

OpenStudy (anonymous):

find the Normals and use the dot product identity on them

OpenStudy (anonymous):

dot product of the normal is -1, what do i do with this?

OpenStudy (anonymous):

A dot B = |A|*|B|*.........?

OpenStudy (anonymous):

how do i get an angle?

OpenStudy (anonymous):

if beta is the angle then\[\cos \beta = n1 \dot n2 /( magnitude of n1) * magnitude of n2)\]

OpenStudy (anonymous):

numerator has dot product

OpenStudy (anonymous):

is n1 and n2 = <-5,-2,4> and <3,3,5>??

OpenStudy (anonymous):

good, you are on right track

OpenStudy (anonymous):

yeah i used those and i got theta = 91.3026 degrees but it says that is wrong?

OpenStudy (anonymous):

what are values of magnitude of n1 and n2 respectively

OpenStudy (anonymous):

magnitude of <-5,-2,4> = sqrt(45) magnitude of <3,3,5> = sqrt(43)

OpenStudy (anonymous):

take the product and what do you have

OpenStudy (anonymous):

43.9886349

OpenStudy (anonymous):

whats in the numerator

OpenStudy (anonymous):

ok i got it correct thanks!!

OpenStudy (anonymous):

great